OpenStudy (spring98):

simplify 5^-4/5^3. Do I add the exponents?

OpenStudy (cggurumanjunath):

5^(-4-3)

OpenStudy (spring98):

so will the answer be 5^7 @CGGURUMANJUNATH

OpenStudy (jhannybean):

\[\large \frac{x^n}{x^m} = x^{n-m}\] therefore \[\frac{5^{-4}}{5^3} = 5^?\]

OpenStudy (spring98):

5^7

OpenStudy (jhannybean):

Not quite. Remember that we are SUBTRACTING the power in the numerator from the power in the denominator.

OpenStudy (spring98):

o so it would be 5^-1 right?

OpenStudy (cggurumanjunath):

5^(-7)

OpenStudy (jhannybean):

no. -4 - 3 = - (4+3) = ?

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^{4} 5^{3}}\]

OpenStudy (spring98):

1/5^7

OpenStudy (jhannybean):

Yes that's another way to do it @yinkim52001

OpenStudy (jhannybean):

Good, and what is 5\(^7\)?

OpenStudy (anonymous):

which is\[\frac{ 1 }{ 5^{7} }\]

OpenStudy (anonymous):

thats your answer

OpenStudy (anonymous):

i know there's another way just show her the steps @Jhannybean

OpenStudy (spring98):

thanks for your help.

OpenStudy (anonymous):

No problem